# multiple directions/.style=for tree=#1, phantom, for relative level=1no edge, delay=!c.content/.pgfmath=content("!u"), before computing xy=l=0,s=0, multiple directions/.default=, grow subtree/.style=for tree=grow=#1, grow’ subtree/.style=for tree=grow’=#1

### v1

@@ 5.2 Uniqueness of the message passing distribution (Theorem 2)
But this is precisely the opposite of $\overline{u}_{\frac{H}{2}}>x$, hence $g(x,w_{i})=0$.
@@
Third, if $x$ is not in the previous two intervals, then let $i_{0}=i_{0}(x,w_{i})$ be the last time the nested maximums evaluate to zero. Then $g$ being one is equivalent to:
@@
and $i_{0}$’s value does not depend on the last values of $w_{i}$, so it does not depend on $w_{H}$ on those intervals (one can see $i_{0}$ as a stopping time on $x$ and $w_{i}$). By condition 4. of the lemma, none of these inequalities change when we modify $w_{H}$ by the prescribed $\varepsilon$, so $g(x,w_{i})$’s value does not change either.
[…]
$\mathbb{P}_{Z_{h_{1}}\sim\zeta_{h_{1}}}(x_{0}\leq Z_{h_{1}}<x_{0}+\varepsilon)=\mathbb{P}_{Z_{h_{2}}\sim\zeta_{h_{2}}}(x_{0}\leq Z_{h_{2}}<x_{0}+\varepsilon).$
This being true for all $x_{0}\in\mathrm{supp}(Z_{h_{1}})$ or $x_{0}\in\mathrm{supp}(Z_{h_{2}})$, this shows that $Z_{1}$ and $Z_{2}$ must be of the same law, hence $h_{1}=h_{2}$.
If $\hat{\pi}_{1}=0$, take the smallest $p\geq 1$ such that $\hat{\pi}_{p}>0$ then condition on a sequence of $p$ children at each step up to depth $H$, by choosing the weights on the other $p-1$ children correctly at each step so that the $\max$ evaluates on the first children, we can recover the previous reasoning as the other branches will not intervene in the recursion. ∎
@@ 6.1 Optimal matching performance and density
In this entire subsection, we reuse the previous notation: $(G_{n})$ is a sequence of random graphs that converges locally to a UBGW of reproduction law $\pi$ and weight law $\omega$ on the edges with $\pi$ and $\omega$ integrable and $\omega$ atomless. Let $\phi$ be the generating function of $\pi$ and $\hat{\phi}$ be $\frac{\phi^{\prime}}{\phi^{\prime}(1)}$. In particular, $\mathbb{E}_{N\sim\pi}[N]=\phi^{\prime}(1)$. The function $h$ is the solution to Equation 9, and the family $Z_{h}(u,v)$ is the associated process on $\mathbb{T}$. Finally, $\zeta_{h}$ the law for which $h$ is the cumulative distribution function.
@@
The asymptotics of the average cost per edge and edge density can be computed on the limiting tree. Let $W$ of law $\omega$, $Z$ and $Z^{\prime}$ of law $\zeta_{h}$ such that $(W,Z_{h},Z_{h}^{\prime})$ are mutually independent. Then:
@@
In particular, this implies the not-so-obvious fact that every solution $h$ to Equation (9) shares the same value for $h(0)$.
@@
The only thing to prove is the last equality $\mathbb{E}\left[\mathbbm{1}_{Z_{h}+Z_{h}^{\prime}<W}\right]=\frac{1-\phi(\hat{\phi}^{-1}(h(0)))}{\phi^{\prime}(1)}$. Let $g(u)=\hat{\phi}(1-u)$, $g^{-1}(u)=1-\hat{\phi}^{-1}(u)$, $G(u)=-\frac{\phi(1-u)}{\phi^{\prime}(1)}$ be a antiderivative of $g$:
@@
$=\int_{\mathbb{R}}\mathbb{E}_{W}[h(W-s)]\mathrm{d}h(s)$
Using the fact that $h$ is solution to Equation 9 we get
@@
From there, elementary calculus gives
@@
To the best of our knowledge, not many nice formulas like the previous one exist. We can however state the following, which generalises Corollary 26.
@@
We can compute the asymptotic of the conditional law of a vertex being matched in $\mathbb{M}_{\mathrm{opt}}(G_{n})$ when its degree is conditioned to be $(k+1)$, for a vertex $v$ chosen uniformly in $G_{n}$:
@@
We will prove the edge-rooted Equation (35). Convergence is obtained by applying Theorem 1 to the local function $f=\mathbbm{1}_{o=(o_{-},o_{+})\in\mathbb{M}\deg(o_{+})=k}$. The function $f$ is $1-$local so we condition on the $1-$neighbourhood.
@@
$N_{-}$ is unrestricted so $Z(o_{+},o_{-})$ is simply of law $\zeta_{h}$, whereas $Z(o_{-},o_{+})=\max_{1\leq i\leq N_{+}}(0,w_{+,i}-Z_{+,i})$. Then the cdf of $Z_{o_{-},o_{+}}$ conditionally on $N_{+}=k$ satisfies:
$h_{N}(t)=\mathbbm{1}_{t\geq 0}\left(1-\mathbb{E}_{W}[h(W-t)]\right)^{k}$
Indeed,
$\mathbb{P}(Z(o_{-},o_{+})\leq t)$ $=\mathbb{P}\left(\max_{1\leq i\leq N_{+}}(0,w_{+,i}-Z_{+,i})\leq t|N_{+}=k\right)$
@@
$=\mathbbm{1}_{t\geq 0}\left(1-\mathbb{E}_{W}[h(W-t)]\right)^{k}.$
Notice that $1-\mathbb{E}_{W}[h(W-t)]=\hat{\phi}^{-1}(h(t))$ by Equation (9). Thus:
$h_{N}(t)=\mathbbm{1}_{t\geq 0}\hat{\phi}^{-1}(h(t))^{k}.$
Now we integrate with respect to the value of $Z_{o_{-},o_{+}}$:
@@
If we push the conditioning further, the expressions depend on the entire shape of the function $h$:
@@
We can compute the asymptotic of the conditional law of a "gap" on one side of an edge of $\mathbb{M}_{\mathrm{opt}}(G_{n})$ when the side of the gap is conditioned to be of degree k, let $h_{w}$ be the cdf of $\omega$, then:
$\lim_{n\rightarrow\infty}\mathbb{P}\left(\text{no children of }v\text{ in }T^{v}\text{ is matched by }\mathbb{M}_{\mathrm{opt}}(G_{n})|\deg(v)=k+1,(u,v)\in\mathbb{M}_{\mathrm{opt}}(G_{n})\right)$
@@
Apply Theorem 1 to the corresponding $2-$local function as we have done before to obtain convergence. Write $v_{i}$ the children of $o_{-}$ in $\mathbb{T}^{o_{-}}$, $w_{i}$ the weight of $(o_{-},v_{i})$, $v_{i,j}$ the children of $v_{i}$ and $w_{i,j}$ the weight of $(v_{i},v_{i,j})$, $Z_{i,j}=Z(v_{i},v_{i,j})$, $Z_{i}=Z(o_{-},v_{i})$, $Z_{o_{+},o_{-}}=Z(o_{+},o_{-})$ and $w_{o}=w(o_{+},o_{-})$. Finally, write $N$ the number of children of $o_{-}$ and $N_{i}$ the number of children of $v_{i}$. See Figure 19 for an illustration. Fix $k\in\mathbb{N}$, the condition rewrites as :
@@
On $B$, we know that $N=k$, but we also know that either $w(v_{i},o_{-})-Z_{(}v_{i},o)<\max_{u\sim v_{i}}(w(v_{i},u)-Z(v_{i},u))$, or both are negative, because else, $v_{i}$ would be matched with $o_{-}$ who is already matched to $o_{+}$. So $A\cap B$ can be rewritten as:
@@
We also notice that the event on the right can be broken down into:
@@
Every variable appearing in this expression are now independent, $w_{i},w_{o}$ of law $\omega$, $N$ of law $\hat{\pi}$, $Z_{i},Z(o_{-},o_{+})$ of law $\zeta_{h}$. So we can compute its probability.
@@ 6.2 Exponential edge weights
To the best of our knowledge, the case where $\omega$ is an exponential law is the only case where $h$ is explicitly computable, and it is no surprise that this case has already been solved by Gamarnik et al. in [gamarnik2003maximum] in the case of Erdös-Renyi graphs and $d-$regular graphs. Indeed let us look at Equation (9) when $W$ is of law $\text{Exp}(1)$.
@@
Now, consider the map $f:x\in[0,1]\mapsto\int_{\mathbb{R}_{+}}\hat{\phi}(1-e^{-u}x)e^{-u}\mathrm{d}u$. We have $f(0)=1,f(1)=\int_{\mathbb{R}_{+}}\hat{\phi}(1-e^{-u})e^{-u}\mathrm{d}u<1$ as long $\hat{\phi}\neq 1$ (in which case the tree is empty), and $f$ is strictly decreasing. Thus, there exists a unique $K$ satisfying the equation and the solution $h$ is unique and explicit, as expected.
6.3 Extensions
To conclude this work, we now give a few possible generalisations. We only give the main ideas and stay purposely light on details.
6.3.1 Multi-type UBGW tree and Stochastic Block Model
Our results should extend to multi-type UBGW trees which also appear in [bordenave2012matchings] but in the unweighted case. These trees appear as local limits for Stochastic Block Models defined as follows. Let $V_{n}=\{1,...,n\}$, and let $k\in\mathbb{N}^{*}$ be the number of types. Let $(\alpha_{1},....,\alpha_{k})$ be strictly positive numbers such that $\sum_{i=1}^{k}\alpha_{k}=1$. For $(i,j)\in\{1,...,k\}^{2}$ fix $c_{i,j}>0$. Partition $V_{n}$ into $V_{n,1}...,V_{n,k}$ such that $\left||V_{n,i}|-\alpha_{i}n\right|<1$, we may have to add one dummy vertex at the end to deal with the fact that $\alpha_{i}n$ may not be an integer, but it doesn’t change the asymptotic behaviour. The (sparse) stochastic block model is the random graph generated on $V_{n}$ so that independently for $x\in V_{n,i}$ and $y\in V_{n,j}$, $\mathbb{P}((x,y)\in E_{n})=p_{i,j}=\frac{c_{i,j}}{n}$.
@@ 6.3.2 Vertex weights
$t$ $\mapsto{}\mathbb{P}_{W\sim\omega}(t\geq W)\hat{\phi}(1-f_{-}(-t))$
with $f_{-}$ the lower continuous version of $f$ and use Schauder’s theorem to recover a solution to the RDE (37).
@@
6.3.3 Maximum subgraph satisfying random capacity constraints
@@
In the deterministic case, when $c(v)\equiv 1$, we recover a matching, when $c(v)\equiv k$, a maximal subgraph of degree less than $k$.
@@
$Z(u,v)=\max\left(0,\underset{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u\end{subarray}}{\overset{[c(v)]}{\max}}\,w(v,u^{\prime})-Z(v,u^{\prime})\right),$
@@
$h(t)=\mathbbm{1}_{t\geq 0}\,\sum_{l=0}^{\infty}c_{l}\sum_{p=0}^{l-1}\frac{\left(\mathbb{E}_{W\sim\omega}\left[h_{-}(W-t)\right]\right)^{p}}{p!}\hat{\phi}^{(p)}(1-\mathbb{E}_{W\sim\omega}\left[h_{-}(W-t)\right])$
@@
$h(t)=\mathbbm{1}_{t\geq 0}\Phi(\mathbb{E}_{W\sim\omega}\left[h_{-}(W-t)\right])$
@@
ENSAE (Fairplay team), France

### v2

@@ 5 Uniqueness of the message passing distribution : Theorem 3
But this is precisely the opposite of $\overline{u}_{\frac{H}{2}}>x$, hence $g(x,w_{0:H})=0$.
@@
Third, if $x$ is not in the previous two intervals, then let $i_{0}=i_{0}(x,w_{0:H})$ be the last time the nested maximums evaluate to zero. Then $g$ being one is equivalent to:
@@
and $i_{0}$’s value does not depend on the last values of $w_{i}$, so it does not depend on $w_{H}$ on those intervals (one can see $i_{0}$ as a stopping time on $x$ and $w_{i}$). By condition 4. of the lemma, none of these inequalities change when we modify $w_{H}$ by the prescribed $\varepsilon$, so $g(x,w_{0:H})$’s value does not change either.
To conclude, the only change in the integral occurs around $x_{0}$, hence for all $\varepsilon$ small enough,
$\mathbb{P}_{Z_{h}\sim\zeta_{h}}(x_{0}\leq Z_{h}<x_{0}+\varepsilon).$
does not depend on $h$. This being true for all $x_{0}\in\mathrm{supp}(Z_{h})$, we have thus shown that the law of $h$ is unique and that the density of $h$ at $x_{0}$ can be recovered by as $\frac{\partial G(w_{0},...,w_{H})}{\partial w_{H}}$ for correctly chosen $w_{0},....,w_{H}$.
If $\hat{\pi}_{1}=0$, take the smallest $p\geq 1$ such that $\hat{\pi}_{p}>0$, then condition on weights such that every path from $o_{+}$ on the $+$ side sees the same weights as in the case $\hat{\pi}_{1}>0$. A little bit of analysis shows that a similar proof holds. ∎
@@ 6.1 Optimal matching performance and density
In this entire subsection, we reuse the previous notation: $(G_{n})$ is a sequence of random graphs that converges locally to a UBGW of reproduction law $\pi$ and weight law $\omega$ on the edges with $\pi$ and $\omega$ integrable and $\omega$ atomless. Let $\phi$ be the generating function of $\pi$ and $\hat{\phi}$ be $\frac{\phi^{\prime}}{\phi^{\prime}(1)}$. In particular, $\mathbb{E}_{N\sim\pi}[N]=\phi^{\prime}(1)$. The function $h$ is the solution to Equation (10), and the family $Z_{h}(u,v)$ is the associated process on $\mathbb{T}$. Finally, $\zeta_{h}$ the law for which $h$ is the cumulative distribution function.
@@
Under the same assumptions as Theorem 2, the asymptotics of the average cost per edge and edge density can be computed on the limiting tree. Let $W$ of law $\omega$, $Z$ and $Z^{\prime}$ of law $\zeta_{h}$ such that $(W,Z_{h},Z_{h}^{\prime})$ are mutually independent. Then:
@@
The only thing to prove is the last equality $\mathbb{E}\left[\mathbbm{1}_{Z_{h}+Z_{h}^{\prime}<W}\right]=\frac{1-\phi(\hat{\phi}^{-1}(h(0)))}{\phi^{\prime}(1)}$. Indeed, the other identities are direct consequences of Theorem 2. Let us start with a simple calculation:
@@
$=\int_{\mathbb{R}}\mathbb{E}_{W}[h(W-s)]\mathrm{d}h(s).$
Using the fact that $h$ is solution to Equation (10) we get
@@
Let $g(u)=\hat{\phi}(1-u)$, $g^{-1}(u)=1-\hat{\phi}^{-1}(u)$, $G(u)=-\frac{\phi(1-u)}{\phi^{\prime}(1)}$ be an antiderivative of $g$. From there, elementary calculus gives
@@
We can also state the following, which generalises Corollary 3.
@@
Under the same assumptions as Theorem 2, the asymptotics of the law of a vertex being matched in $\mathbb{M}_{\mathrm{opt}}(G_{n})$ when its degree is conditioned to be $(k+1)$, for a vertex $v$ chosen uniformly in $G_{n}$, is given by :
@@
We will prove the edge-rooted Equation (41). Convergence is obtained by applying Theorem 2 to the local function $f=\mathbbm{1}_{o=(o_{-},o_{+})\in\mathbb{M},\deg(o_{+})=k}$. The function $f$ is $1-$local so we condition on the $1-$neighbourhood.
@@
$N_{-}$ is unrestricted so $Z(o_{+},o_{-})$ is simply of law $\zeta_{h}$, whereas $Z(o_{-},o_{+})=\max_{1\leq i\leq N_{+}}(0,w_{+,i}-Z_{+,i})$. This leads to computing
$\mathbb{P}\left(Z(o_{-},o_{+}\right)\leq t|N_{+}=k)$ $=\mathbb{P}\left(\max_{1\leq i\leq N_{+}}(0,w_{+,i}-Z_{+,i})\leq t|N_{+}=k\right)$
@@
$=\mathbbm{1}_{t\geq 0}\left(1-\mathbb{E}_{W}[h(W-t)]\right)^{k}$
$=\mathbbm{1}_{t\geq 0}\hat{\phi}^{-1}(h(t))^{k}$
by noticing that $1-\mathbb{E}_{W}[h(W-t)]=\hat{\phi}^{-1}(h(t))$ according to Equation (10). Define:
$h_{k}(t)=\mathbbm{1}_{t\geq 0}\hat{\phi}^{-1}(h(t))^{k}.$
We have just seen that $h_{k}$ is the cdf of $Z_{o_{-},o_{+}}$ conditionally on $N_{+}=k$ Now we integrate with respect to the value of $Z_{o_{-},o_{+}}$:
@@
We did not manage to find further nice formulas that only depend on $h(0)$. Indeed, if we push the conditioning further, the expressions depend on the entire shape of the function $h$. This is the case for instance for the probability of presence of a gap in the matching:
@@
Under the same assumptions of Theorem 2, we can compute the asymptotics of the law of a "gap" on one side of an edge of $\mathbb{M}_{\mathrm{opt}}(G_{n})$ when the side of the gap is conditioned to be of degree k. Let $h_{w}$ be the cdf of $\omega$, then:
$\lim_{n\rightarrow\infty}\mathbb{P}\left(\text{no children of }v\text{ in }T_{(u,v)}\text{ is matched by }\mathbb{M}_{\mathrm{opt}}(G_{n})|\deg(v)=k+1,(u,v)\in\mathbb{M}_{\mathrm{opt}}(G_{n})\right)$
@@
As in the previous corollaries, we apply Theorem 2 to the corresponding $2-$local indicator function of the event to obtain the convergence to the corresponding probability on the tree. Write $v_{i}$ the children of $o_{-}$ in $\mathbb{T}_{(o_{+},o_{-})}$, $w_{i}$ the weight of $(o_{-},v_{i})$, $v_{i,j}$ the children of $v_{i}$ and $w_{i,j}$ the weight of $(v_{i},v_{i,j})$, $Z_{i,j}=Z(v_{i},v_{i,j})$, $Z_{i}=Z(o_{-},v_{i})$, $Z_{o_{+},o_{-}}=Z(o_{+},o_{-})$ and $w_{o}=w(o_{+},o_{-})$. Finally, write $N$ the number of children of $o_{-}$ and $N_{i}$ the number of children of $v_{i}$. See Figure 19 for an illustration. Fix $k\in\mathbb{N}$, the condition rewrites as:
@@
On $B$, we know that $N=k$, but we also know that either $w(v_{i},o_{-})-Z(v_{i},o)<\max_{u\sim v_{i}}(w(v_{i},u)-Z(v_{i},u))$, or both are negative, because else, $v_{i}$ would be matched with $o_{-}$ who is already matched to $o_{+}$. So $A\cap B$ can be rewritten as:
@@
We also notice that the event on the right can be written as
@@
Every variable appearing in this expression are now independent, namely, $w_{i},w_{o}$ of law $\omega$, $N$ of law $\hat{\pi}$, $Z_{i},Z(o_{-},o_{+})$ of law $\zeta_{h}$. So we can compute its probability.
@@ 6.2 Exponential edge weights
The case where $\omega$ is an exponential law is one of the only cases where $h$ is explicitly computable, and it has been solved by Gamarnik et al. in [gamarnik2003maximum] in the case of Erdös-Renyi graphs and $d-$regular graphs. Our method generalizes their result to a general reproduction distribution as follows. Let us look at Equation (10) when $W$ is of law $\text{Exp}(1)$.
@@
Now, consider the map $f:x\in[0,1]\mapsto\int_{\mathbb{R}_{+}}\hat{\phi}(1-e^{-u}x)e^{-u}\mathrm{d}u$. We have $f(0)=1,f(1)=\int_{\mathbb{R}_{+}}\hat{\phi}(1-e^{-u})e^{-u}\mathrm{d}u<1$ as long $\hat{\phi}\neq 1$ (in which case the tree is empty), and $f$ is continuous and strictly decreasing. Thus, there exists a unique $K$ satisfying the equation and the solution $h$ is unique and explicit, as expected.
7 Extensions
To conclude, we now give a few possible generalisations. We only give the main ideas and stay purposely light on details.
7.1 Multi-type UBGW tree and Stochastic Block Model
Our results extends to multi-type UBGW trees which also appear in [bordenave2012matchings] but in the unweighted case. These trees appear as local limits for Stochastic Block Models defined as follows. Let $V_{n}=\{1,...,n\}$, and let $k\in\mathbb{N}^{\star}$ be the number of types. Let $(\alpha_{1},....,\alpha_{k})$ be strictly positive numbers such that $\sum_{i=1}^{k}\alpha_{k}=1$. For $(i,j)\in\{1,...,k\}^{2}$ fix $c_{i,j}>0$. Partition $V_{n}$ into $V_{n,1}...,V_{n,k}$ such that $\left||V_{n,i}|-\alpha_{i}n\right|<1$, we may have to add one dummy vertex at the end to deal with the fact that $\alpha_{i}n$ may not be an integer, but it doesn’t change the asymptotic behaviour. The (sparse) stochastic block model is the random graph generated on $V_{n}$ so that independently for $x\in V_{n,i}$ and $y\in V_{n,j}$, $\mathbb{P}((x,y)\in E_{n})=p_{i,j}=\frac{c_{i,j}}{n}$.
@@ 7.2 Vertex weights
$t$ $\mapsto{}\mathbb{P}_{W\sim\omega}(t\geq W)\hat{\phi}(1-f(-t))$
with $f$ the lower continuous version of $f$ and use Schauder’s theorem to recover a solution to the RDE (43).
@@
7.3 Maximum subgraph satisfying random capacity constraints
@@
In the deterministic case, when $c(v)\equiv 1$, we recover a matching, when $c(v)\equiv k$, a maximal subgraph with degrees less than or equal to $k$.
@@
$Z(u,v)=\max\left(0,\underset{\begin{subarray}{c}u^{\prime}\sim v\\ u^{\prime}\neq u\end{subarray}}{\overset{[c(v)]}{\max}}\,(w(v,u^{\prime})-Z(v,u^{\prime}))\right),$
@@
$h(t)=\mathbbm{1}_{t\geq 0}\,\sum_{l=0}^{\infty}c_{l}\sum_{p=0}^{l-1}\frac{\left(\mathbb{E}_{W\sim\omega}\left[h(W-t)\right]\right)^{p}}{p!}\hat{\phi}^{(p)}(1-\mathbb{E}_{W\sim\omega}\left[h(W-t)\right])$
@@
$h(t)=\mathbbm{1}_{t\geq 0}\Phi(\mathbb{E}_{W\sim\omega}\left[h(W-t)\right])$
@@
GENES, Fairplay joint team, CREST France